# Mechanical hysterons with tunable interactions of general sign

### v2

@@ Abstract
The non-reciprocity of the interaction can now immediately be seen, as $J_{ij}$ depends on the angular interval $2\Delta\theta_{j}=\theta_{j}^{+}-\theta_{j}^{-}$ but not on $\theta_{i}$. The larger splitting of hysteron 2 in Fig. 2A is thus due to the larger angular interval of hysteron 1. To test this result quantitatively, Fig. 2C compares Eq. 5 with the value of $\gamma_{i}^{\pm}(+)-\gamma_{i}^{\pm}(-)$ from the full kinematic model and the experiments. The comparison shows that Eq. 5 captures the observed non-reciprocity, and that its dependence on $\ell$ is also in good agreement.
To test the small-$\theta$ model further, Fig. 3 compares the kinematic model with the small-$\theta$ model for a pair of hysterons with parallel springs (Fig. 3A) and crossed springs (Fig. 3B). The parameters are otherwise the same as in Fig. 2. The small-$\theta$ model captures the gross trends of all 8 switching thresholds in both configurations. In particular, the shift in $[\gamma_{i}^{\pm}(+)+\gamma_{i}^{\pm}(-)]/2$ is captured by the small-$\theta$ model, and is due to the last term in Eq. 4, which measures the difference between the mean rotor angle $\overline{\theta_{j}}$ and the current rotor angle $\theta_{i}$ [recall that the ($+$) transition is out of the (--) state and vice versa]. This shift is the most apparent in the thresholds for hysteron 1 in Fig. 3, as that rotor is bounded by larger angles. In general, the pair $\gamma_{i,0}^{+}(\pm)$ and the pair $\gamma_{i,0}^{-}(\pm)$ shift in opposite directions, with the sign of the shift set by $g_{ij}$.
Finally, we note that the thresholds and interactions in Eqs. 4 and 5 are independent of $w$. This means that one can couple together hysterons that are not adjacent, while maintaining the same design space for interactions. We also note that Eqs. 4 and 5 are readily adapted to heterogeneous $L=L_{i}$ or $\ell=\ell_{ij}$, which can be used to tune the coupling to the global drive or the relative pairwise interactions in larger systems. We make use of these features later on.
Figure 4: Suite of behaviors for an interacting hysteron pair. A. Transition graphs obtained in the experiment, showing all states and transitions accessible from the initial state, ---- (Supplementary Movies 1-4). Graphs $i$ and $ii$ are “Preisach” graphs that do not require interactions [15], whereas graphs $iii$-$vi$ require cooperative interactions. To show how one can move between different transition graphs, we demonstrate them all using cooperative interactions, without changing the driving or coupling springs (“Configuration 1” in C). Graphs $vii$-$ix$ require frustrated interactions; we obtain them with a second set of driving and coupling springs (“Configuration 2” in C). B. Switching thresholds calculated from the model, using the experimental parameters. C. Experimental parameters. We fix $x_{1}=0$ cm, $y=14.5$ cm, $L=6.3$ cm, $\ell=4.7$ cm, and $w=15.2$ cm. D. Schematic illustrating the role of interactions on hysteron 2 in graphs $i$-$ix$. Top row: hysteron 1 is (+)(+)( + ); bottom row: hysteron 1 is ()(-)( - ). Smaller cooperative interactions are required to create graphs $i$-$vi$ in B, whereas larger frustrated interactions are needed for graphs $vii$-$ix$.
Targeting specific transition graphs.— The above results suggest that our mechanical hysterons should be able to produce many of the behaviors of the general hysteron model of Eq. 1. To develop a better sense of this presumed generality, and to see whether there are experimental limitations that hamper it, we begin by targeting different transition graphs for two coupled hysterons. As we do so, we also investigate a parallel question: To what extent can different functionalities be accessed while changing only a small number of system parameters?
To this end, we start by demonstrating four distinct transition graphs in a sequence where we alter only the mounting position of driving spring 2, thereby changing only $x_{2}$. Supplementary Movie 1 shows this sequence of behaviors, which correspond to transition graphs $i$-$iv$ in Fig. 4A. The experimental parameters are tabulated in Fig. 4C.
To understand how the system moves between these four transition graphs, we plug these parameters into the kinematic model, calculate the switching thresholds, and plot them in Fig. 4B. The switching thresholds of hysteron 1 are constant throughout this sequence, as they depend on the state of hysteron 2 but not its driving. The thresholds for hysteron 2 move in concert without changing their spacing, as changing $x_{2}$ shifts each threshold by $x_{2}$. Careful inspection of the switching thresholds in Fig. 4B shows that they indeed produce graphs $i$-$iv$ in Fig. 4A, faithful to the experiment.
Next, we access transition graph $v$ by decreasing $\Delta\theta_{1}$, increasing $\Delta\theta_{2}$, and changing $x_{2}$ (Supplementary Movie 2). We access transition graph $vi$ by setting $\Delta\theta_{1}$ = $\Delta\theta_{2}$ and changing $x_{2}$ once again (Supplementary Movie 3). This exhausts all possible transition graphs for two hysterons with cooperative interactions [15].
Three additional transition graphs can be obtained by using frustrated interactions [15]; they are graphs $vii$-$ix$ in Fig. 4A. Supplementary Movie 4 shows how we can realize these graphs in the experiment while varying only $x_{2}$ to navigate between them. All three graphs feature at least one “horizontal avalanches”. To understand how they occur, start in the --- - state in graph $vii$. The corresponding thresholds in Fig. 4B show that upon increasing $\gamma$, hysteron 2 switches first, as $\gamma_{2}^{+}(-)<\gamma_{1}^{+}(-)$. Now in the +-+- + state, the next event is at $\gamma=\gamma_{1}^{+}(+)$ into +++++ +. But this state is not stable, as this $\gamma$ is below $\gamma_{2}^{-}(+)$, which brings the system immediately to ++-+ -. An analogous process underlies each of the ++-$\leftrightarrow$+-- + ↔ + - avalanches in graphs $vii$-$ix$.
Among these graphs, perhaps the most exotic is $ix$, which navigates the sequence: --- -, +-+- +, ++-+ -, +++++ + under unidirectional driving. This behavior constitutes a 2-bit analog-to-digital converter [15]. We now describe the process used to target this transition graph, to illustrate how behaviors may be rationally designed in our system. First, we need a strong coupling between the global drive and hysteron 1, so that it is ()(-)( - ) at small $\gamma$ and (+)(+)( + ) at large $\gamma$. We also need a strong frustrated interaction between hysterons 1 and 2 to navigate the ++-$\leftrightarrow$+-- + ↔ + - avalanches. In particular, the interaction of hysteron 1 on hysteron 2 must open up a central interval, $\gamma_{2}^{+}(-)<\gamma<\gamma_{2}^{-}(+)$ where the state of hysteron 2 depends only on the current state of hysteron 1. Figure 4D illustrates this condition, and an examination of graphs $vii$-$ix$ in Fig. 4 shows that they satisfy it. We target this condition by selecting a large $k_{1,2}$ and a sufficiently large $\theta_{1}^{\pm}$; we find that the same ±$\pm 38.8^{\circ}$ used in the other designs works here too. Finally, a weak coupling of hysteron 2 to the global drive enables a transition into the +++++ + state at large $\gamma$ and into --- - at small $\gamma$. These considerations lead to a tentative design for the spring constants: $k_{1}\approx k_{1,2}\gg k_{2}$. Here we use $k_{1}=k_{1,2}=0.46$ N/cm and $k_{2}=0.07$ N/cm, which we found to suffice.
Two additional transition graphs are possible for “mixed” interactions, where the signs of $J_{ij}$ and $J_{ji}$ differ [15]. We do not see a way to realize such interactions in our system.
Figure 5: Targeting a latching behavior. A. Coupled mechanical hysterons that are stable in the ++-+ - and --- - states at $\gamma=0$. Driving up to $\gamma=10$ cm and back to 00 sets the system in ++-+ -. Driving up to $\gamma=32$ cm and back to 00 resets the system to --- -. Geometric parameters: $y=9.4$ cm, $L=6.3$ cm, $\ell=4.7$ cm, $w=15.2$ cm, $x_{1}=-8.1$ cm, $x_{2}=-16.4$ cm, $\theta_{1}^{\pm}=\pm 18.3^{\circ}$, and $\theta_{2}^{\pm}=\pm 38.8^{\circ}$. All springs have $k=0.19$ N/cm and $x_{0}=4.1$ cm. B. Switching thresholds in the experiment and the model are in good agreement and follow the requisite ordering for latching [24].
Latching.— The ability to manifest non-reciprocal frustrated interactions in this system allows one to access a behavior termed “latching”. This behavior was proposed by Lindeman et al. [24] to explain how amorphous solids form multiple memories of asymmetric driving. Here we demonstrate it with our mechanical hysterons in Fig. 5A and Supplementary Movie 5. Starting in the --- - state at $\gamma=0$, we drive the system to $\gamma=10$ cm, which puts the hysterons in the ++-+ - state. This state remains stable as we return $\gamma$ back to 00; hysteron 1 is now “latched” in the (+)(+)( + ) state. A larger driving amplitude to $\gamma=32$ cm and back to 00 releases hysteron 1 back to ()(-)( - ) via a sequence of intermediate states: ++++,+$\rightarrow$+$\rightarrow$-+$\rightarrow$--+ - → + + → - + , → - -. This behavior necessarily violates return-point memory [29, 30, 26], which would stipulate that the same state would recur each time the driving returns to its common endpoint, $\gamma=0$.
Lindeman et al. [24] showed how latching is tied to the ordering of the switching thresholds. They found that latching requires:
$\gamma_{1}^{-}(-)<0<\gamma_{2}^{-}(-)<\gamma_{2}^{-}(+)<\gamma_{1}^{-}(+)\ ,$ (6a)
$\gamma_{1}^{+}(-)<\gamma_{2}^{+}(-)\ .$ (6b)
We now show how we translate this ordering into a concrete design in our mechanical system. We start by noting that Eq. 6a implies: $0<\gamma_{2}^{-}(+)-\gamma_{2}^{-}(-)<\gamma_{1}^{-}(+)-\gamma_{1}^{-}(-)$, which we rewrite as: $0<-J_{21}<-J_{12}$. We obtain this non-reciprocal frustrated interaction by confining hysteron 1 to a smaller angular interval than hysteron 2 and coupling them together with crossed springs following Eq. 5. Next, we note that adjusting $x_{2}$ moves all thresholds for hysteron 2 in concert without altering their spacing. We thus move $x_{2}$ until $\gamma_{2}^{-}(+)<\gamma_{1}^{-}(+)$ and $\gamma_{1}^{-}(-)<\gamma_{2}^{-}(-)$. Then, we fulfill $\gamma_{1}^{-}(-)<0<\gamma_{2}^{-}(-)$ by moving $x_{1}$ and $x_{2}$ together by equal amounts until those thresholds straddle $\gamma=0$. Finally, $\gamma_{1}^{+}(-)<\gamma_{2}^{+}(-)$ turns out to be satisfied already through the larger angular interval of hysteron 2.
Figure 5B shows the switching thresholds as measured in our experimental realization of latching, using the parameters in the caption. The ordering of thresholds indeed follow Eq. 6. Plugging the experimental parameters into the kinematic model, we obtain the second set of bars, which are in good agreement with the experiment and follow the same ordering.
Our mechanical system allows one to latch and release the hysterons in another way, shown in Supplementary Movie 6. Holding the driving rod fixed at $\gamma=0$, the left rotor can be flipped between ()(-)( - ) the (+)(+)( + ) by hand, and it is stable in either state when the right rotor is ()(-)( - ). From the state ++-+ -, flipping the right rotor to (+)(+)( + ) immediately releases the left rotor to ()(-)( - ), returning the system to --- - upon letting go.
Counting modulo 2.— More sophisticated functions can be realized with more hysterons. Here we target a multiperiodic response that was obtained for simulated hysterons [13, 14, 16]. We start with a design from one of those studies [13], where four hysterons are coupled as in Fig. 6A. When cyclic driving of unit amplitude is applied [31], these parameters produce the steady state drawn with black and red arrows in Fig. 6B, which we determine using the open-source hysteron software package [31]. The sequence of states takes two driving cycles to traverse. These interacting hysterons thus perform a basic computation: discriminating between even and odd numbers of driving cycles, i.e., counting modulo 2.
Figure 6: Counting modulo 2 with four mechanical hysterons. A. Thresholds and interactions that yield a limit cycle with twice the period of the driving, for unit-amplitude cyclic drive [13]. B. Corresponding pathway, where we have colored the two states at the minimum of the driving. Lighter arrows show all other transitions available to the system under arbitrary driving and are not used in the period-2 cycle. States without arrows to them are unstable at all $\gamma$. C. Realizing this pathway with mechanical hysterons. The rotors from left to right correspond to the hysterons from top to bottom in A. The springs connecting rotors 1 and 3 have $k=0.07$ N/cm, and their original $x_{0}=2.5$ cm are extended by steel wire. All other springs have $k=0.19$ N/cm and $x_{0}=4.1$ cm. D. States and transitions in the experiment, which achieves the targeted response. Horizontal segments denote avalanches.
Investigating this pathway in Fig. 6B further, we can see that it features two avalanches (+-+- ++-+- + $\rightarrow$ +++++ ++++++ + and +++++ $+$-+ - $\rightarrow$ --- -++-+ -), as well as a pair of “scrambled” transitions: +++++ ++++++ + $\rightarrow$ +++++ $+$-+ - and --- -+++++ + $\rightarrow$ --- -+-+- +, meaning that they are not consistent with a state-independent ordering of the switching fields [19, 15]. To see this, note that these transitions imply that $\gamma_{3}^{-}>\gamma_{4}^{-}$ when $s_{1}=s_{2}=-1$, and that $\gamma_{3}^{-}<\gamma_{4}^{-}$ when $s_{1}=s_{2}=1$. The presence of either scrambling or avalanches is known to be necessary, but not sufficient, for a multiperiodic response [15].
@@
To reproduce this complex response in the lab, we start by translating the sign and approximate strength of the interactions into a choice of springs and their mounting points on the rotors, which controls the magnitude of the interaction via the product $k_{ij}\ell_{ij}^{2}$. One pair of springs spans a longer distance from rotor 1 to rotor 3; we attach a length of steel wire to extend the reach of these springs while staying within their linear response. We then set the $\Delta\theta_{i}$ for each hysteron to reflect its relative amount of hysteresis in the design. To ensure a repeatable minimum and maximum for the driving, we mount the two linear bearings that hold the driving rod at locations where they will collide with the outermost clamps that support springs 1 and 4, and we use these contacts as the turning points.
@@
The mounting positions $x_{i}$, the angular intervals $2\Delta\theta_{i}$, the coupling positions $\ell_{ij}$, and the two endpoints of the driving offer a total of 14 continuous degrees of freedom. On top of this, we also allow ourselves to mount some of the springs within a pair at unequal $\ell_{ij}$. To target the sequence of states in Fig. 6B, we drive the system to observe its behavior and then adjust one or more of these degrees of freedom to try to obtain more of the desired sequence. This iterative process eventually produced the configuration shown in Fig. 6C; its behavior under cyclic driving is shown in Supplementary Movie 7. To show the resulting steady state, we measure the value of $\gamma$ where each transition occurs, and we plot the sequence of states as a function of time in Fig. 6D. The sequence matches that of Fig. 6B, it takes two periods to traverse, and it repeats indefinitely, as desired.
There is one subtle difference in the pathways: The experiment navigates the downward avalanche via the unstable intermediate state +-+- $+$-+ -, whereas the design in Fig. 6A passes through ++-+ -++-+ - instead. We found that we could match the experiment by changing the bare threshold $\gamma_{1,0}^{-}$ for the top hysteron in Fig. 6A from 00 to $0.85$.
Counting to $n$.— As a final demonstration of computing with mechanical hysterons, Fig. 7 and Supplementary Movie 8 show an accumulator, which counts the number of driving cycles applied to it. A mechanical metamaterial was recently designed with this function [32], using a train of carefully designed elastic beams that make and break contact under cyclic compression. Our design is based on an antiferromagnetic chain, here a series of hysterons with uniform frustrated interactions. We stagger the points where the springs attach to the driving rod, which alternatingly raises and lowers the values of $\gamma^{+}$ and $\gamma^{-}$ for each hysteron. In particular, we set $x_{2m-1}=-L\sin(\theta_{0})$, $x_{2m}=L\sin(\theta_{0})$, and $x_{1}=-3L\sin(\theta_{0})$, where $\theta_{0}=\theta_{i}^{+}=-\theta_{i}^{-}$, although other values can also work. This design breaks the degeneracy of the two ground states so that ++-+ -++-+ -++-+ - has a higher energy than +-+- ++-+- ++-+- +.
Starting in the higher-energy ground state, cycling the system at constant amplitude drives a domain wall down the chain, irreversibly flipping one hysteron each half cycle. These transitions occur because a hysteron at the domain wall experiences two interactions that cancel, allowing it to change state upon the influence of the global drive. The location of the domain wall thus encodes the number of driving cycles; $2n$ hysterons can record this number up to $n$. Once the domain wall reaches the end of the chain, the configuration is stable to further driving.
Figure 7: Counting driving cycles with a chain of mechanical hysterons. Time series showing a domain wall ratcheting down an antiferromagnetic chain, irreversibly flipping one hysteron each half cycle. The system is initialized in the state ++-+ -++-+ -++-+ - and a cycle of driving follows $\gamma:\gamma_{0}\rightarrow-\gamma_{0}\rightarrow\gamma_{0}$. The growing +-+- ++-+- +... phase is highlighted in blue. For ease of construction, the radial bearings are replaced by optical posts in post holders lubricated with silicone oil. Coupling springs: $k=0.17$ N/cm, $x_{0}=4.0$ cm. Driving springs: $k=0.19$ N/cm, $x_{0}=4.1$ cm.
The basic idea of this design can be verified on a computer [31] by applying cyclic driving of any fixed amplitude $\gamma_{0}$, setting thresholds $\gamma_{2m-1}^{+}>-\gamma_{2m-1}^{-}=\gamma_{0}/2$ for the odd hysterons and $\gamma^{\pm}_{2m}=-\gamma^{\mp}_{2m-1}$ for the even hysterons (where $m=1,2,3...$), and setting nearest-neighbor interactions of strength $J_{i,i+1}=-\gamma_{0}$. The thresholds for hysteron 1 should be raised by $\gamma_{0}$ to allow a domain wall to nucleate there.
More precisely, our accumulator design detects cycles of intermediate amplitude. A small-amplitude cycle that stays within the interval $[\max(\gamma_{i}^{-}),\min(\gamma_{i}^{+})]$ will not change the state, and a large-amplitude cycle that overpowers the frustrated interactions can send the system to its final absorbing state. In our idealized example, the system counts cycles of any amplitude between $\gamma_{0}/2$ and $25\gamma_{0}/$ 2.
@@
Discussion.— Here we have demonstrated a design for mechanical hysterons with continuously tunable thresholds and interactions, whose states can be controlled manually or by coupling to an external mechanical drive. Recent work has demonstrated other mechanical designs that can count modulo two [33], count driving cycles up to $n$ [32], or encode an analog mechanical input into a binary string [34]. Although those systems use bistable mechanical elements, each one uses a different strategy to navigate the pathways between its states. In contrast, we have presented a general platform that can be reconfigured to perform all of the above functions, plus a latching behavior that was recently identified in hysteron simulations [24]. Crucial to this versatility was the realization of cooperative and frustrated interactions that can be reciprocal or non-reciprocal.
@@
One principle of our approach was to use nonlinearities sparingly, so that the behaviors of the system could be rationalized as simply as possible. We used linear springs and nearly-frictionless bearings in the experiment, and our analysis using small rotor angles and an affine relationship between torque and the global drive was able to capture the essential behaviors of two interacting hysterons (Fig. 3). Other designs for interacting mechanical hysterons are also being proposed; Shohat and van Hecke [35] used theory and simulations to explore two-dimensional networks of bilinear hysterons as a route to exotic pathways, including mixed interactions and multiperiodic cycles. Understanding what are the general constraints on $\gamma_{i,0}^{\pm}$ and $J_{ij}$ for any realizable metamaterial is an open question.
Designing the response of larger collections of mechanical hysterons presents another set of challenges. Although we report an exact kinematic model and a mapping to switching thresholds and interactions for small rotor angles, in practice we worked at relatively large rotor angles and proceeded through a combination of physical reasoning and trial-and-error. This approach might not work for larger hysteron machines. As an alternative, our kinematic model can provide the basis for computational inverse design methods, which could generate sets of parameters for a particular desired behavior. Another possibility for targeting behaviors in large systems is through supervised learning [36], which was recently implemented in experiments on spring networks [37]. Our system offers a wealth of possible learning degrees of freedom, as there are numerous geometric parameters that can be tuned continuously to affect the switching thresholds, including $x_{i}$, $\theta_{i}^{\pm}$, $\ell_{ij}$, and $L_{i}$. Understanding whether this approach can be used achieve a specific transition graph or a desired response under cyclic drive could be a fertile direction for future work.
The focus of this work has been developing a mechanical hysteron with general interactions (Figs. 1-3), understanding these interactions (Eqs. 3-5), and probing their versatility through a series of targeted behaviors (Figs. 4-7). Some of these behaviors are readily interpreted as basic computations on mechanical inputs, like the accumulator that counts cycles of intermediate amplitude (Fig. 7). Yet, such computation is not limited to a small set of particular designs – rather, it is a general feature of collections of interacting hysterons. This can be realized by interpreting the transition graphs as finite-state machines (FSMs), through the selection of an initial state and an accept state and the choice of an alphabet of driving protocols, as described recently by Liu et al. [21]. For instance, the latching mechanism has a natural two-character alphabet of $\gamma:0\rightarrow 10\rightarrow 0$ as S (set) and $\gamma:0\rightarrow 32\rightarrow 0$ as R (reset). If the initial and accepting states are both --- -, this FSM accepts the empty string and any string ending in R. The computational power of FSMs, paired with the existence of general mechanical hysterons that we have established here, suggests that there is a large, untapped potential for designed materials that can sense, compute, and respond to their mechanical environment.
Acknowledgements.
I thank Nathan Keim for suggesting how to apply a global drive, for providing the parameters used in Fig. 6A, and for many fruitful discussions. I thank Vidyesh Anisetti, Eadin Block, Christian Santangelo, and Jennifer Schwarz for contributing to early hysteron designs, and I thank Martin van Hecke, Nathan Keim, Chloe Lindeman, and Dor Shohat for comments on the manuscript. This work was initiated at the Aspen Center for Physics, which is supported by National Science Foundation grant PHY-2210452.
Appendix A Appendix
@@
Rotor hysteron construction.— We construct each rotor hysterons from two $1/2$”-diameter optical posts attached in a T configuration, which we mount in a radial steel ball bearing affixed to an optical table. We remove the bearing grease to further lower their friction to rotation. We use steel extension springs to couple the rotors to one another and to a 1/2”-diameter steel shaft that is mounted in two linear bearings. This 72”-long rod translates freely along the horizontal direction and provides the global mechanical drive, $\gamma$. We define $\gamma$ as the horizontal displacement of this rod in centimeters, which we measure with a precision of $0.1$ cm via a ruler affixed under the rod.

### v3

@@ Abstract
Non-reciprocal interactions. Returning first to Fig. 2A, a closer study reveals that those interactions are non-reciprocal: The splitting of the blue curves, $\gamma_{2}^{\pm}(-)-\gamma_{2}^{\pm}(+)$, is larger than the splitting of the black curves, $\gamma_{1}^{\pm}(-)-\gamma_{1}^{\pm}(+)$. The non-reciprocity of the interaction can immediately be seen via Eq. 6, where $J_{ij}$ depends on the angular interval $\theta_{j}^{+}-\theta_{j}^{-}$ but does not depend on $\theta_{i}^{\pm}$. The larger splitting of hysteron 2 in Fig. 2A is thus due to the larger angular interval of hysteron 1. To test this result quantitatively, Fig. 2C compares Eq. 6 with the value of $\gamma_{i}^{\pm}(-)-\gamma_{i}^{\pm}(+)$ from the full kinematic model and the experiments. The comparison shows that Eq. 6 captures the observed non-reciprocity, as well as the predicted $\ell^{2}$ scaling.
Having established a physical picture for the switching thresholds and interactions for pairs of mechanical hysterons, the remainder of this article is devoted to targeting more complex behaviors with two, four, and many coupled hysterons.
Hysteron latching. Holding a blueprint for frustrated non-reciprocal interactions allows one to target an exotic behavior termed “latching”. This behavior was proposed by Lindeman et al. Lindeman et al. (2025) to explain how amorphous solids form multiple memories of asymmetric driving. Here we demonstrate it with our mechanical hysterons in Fig. 3A and Supplementary Movie 6.
Starting in the -- state at $\gamma=0$, we drive the system to $\gamma=10$ cm, which puts the hysterons in the ++- state. This state remains stable as we return $\gamma$ back to 0; hysteron 1 is now “latched” in the $(+)$ state. A larger driving amplitude to $\gamma=32$ cm and back to 0 releases hysteron 1 back to ()(-) via a sequence of intermediate states: $-\rightarrow++\rightarrow-+$$\rightarrow--$.
Latching necessarily violates return-point memory Barker et al. (1983); Sethna et al. (1993); Keim et al. (2019), which would stipulate that the same state would recur each time the driving returns to its common endpoint, $\gamma=0$. Indeed, Lindeman et al. Lindeman et al. (2025) showed that latching requires a strong, frustrated, non-reciprocal interaction: $J_{12}<J_{21}<0$. Figure 3B shows that our setup satisfies this requirement. We obtain this frustrated non-reciprocal interaction by confining hysteron 1 to a smaller angular interval than hysteron 2 and coupling them with crossed springs. The full requirements for latching, and how we target them using our mechanical hysterons, is provided in the Supplementary Information.
Counting driving cycles. Frustrated interactions enable other computational behaviors, such as counting the number of mechanical driving cycles that have been applied. Our design uses an antiferromagnetic chain of hysterons with uniform frustrated interactions and uniform hysteresis, $\gamma^{+}_{i}-\gamma^{-}_{i}$. We alternately raise and lower both switching thresholds, $\gamma_{i}^{+}$ and $\gamma_{i}^{-}$, for each hysteron (Fig. 4A). This staggering breaks the degeneracy of the two ground states so that ++-++-... has a higher energy than +-++-+.... Consider now a stable state with one “domain wall” in the bulk where the two ground states meet. Both hysterons at the domain wall experience two interactions that cancel, lowering their barrier to change state via the global drive. This grows the ++-++-... domain by one site for each half-cycle of driving. The location of the domain wall thus encodes the number of driving cycles; $2n$ hysterons can record this number up to $n$.
This design is easily translated to our mechanical system by forming a chain of rotors that are coupled to their nearest neighbors with crossed springs, and attaching them to the driving rod with $x_{i}$ that alternate between negative and positive values (Fig. 4B). In particular, we set $x_{2m-1}=-L\sin(\theta_{0})$, $x_{2m}=L\sin(\theta_{0})$, where $\theta_{0}=\theta_{i}^{+}=-\theta_{i}^{-}$, although other values can also work (see the Supplementary Information). Finally, we set $x_{1}=-3L\sin(\theta_{0})$ at the start of the chain to nucleate a domain wall there. Supplementary Movie 7 and Fig. 4B show the system evolution, starting from the higher-energy ground state. Once the domain wall reaches the end of the chain, the configuration is stable to further driving.
Counting modulo 2. Figure 5A shows a system of four hysterons that, when driven at unit amplitude, produces a steady state that takes two cycles to traverse. It concludes even cycles in the state --+-+, and it concludes odd cycles in the state --++- (Fig. 5B). The system thus performs a basic computation: discriminating between even and odd numbers of driving cycles, i.e., counting modulo 2. Such a multiperiodic response has been identified in simulated hysteron systems Keim and Paulsen (2021); Lindeman and Nagel (2021); Szulc et al. (2022). Here we realize it in the lab.
@@
We start by translating the sign and relative strength of the interactions in Fig. 5A into a choice of springs and their mounting points on the rotors. One pair of springs spans a longer distance from rotor 1 to rotor 3; we attach a length of steel wire to extend these springs while staying within their linear response. We then set $\theta_{i}^{+}-\theta_{i}^{-}$ for each hysteron to reflect its hysteresis in the design. To ensure constant-amplitude driving, we use the linear bearings that hold the driving rod as a “sandwich” that sets the endpoints of the driving.
@@
The mounting positions $x_{i}$, the angular intervals $\theta_{i}^{+}-\theta_{i}^{-}$, the coupling positions $\ell_{ij}$, and the two endpoints of the driving offer a total of 14 continuous degrees of freedom. We also allow ourselves to mount some of the springs within a pair at unequal $\ell_{ij}$. To target the sequence of states in Fig. 5B, we drive the system to observe its behavior and then adjust one or more of these degrees of freedom to try to obtain more of the desired sequence. This iterative process eventually produced the configuration shown in Fig. 5C; its behavior under cyclic driving is shown in Supplementary Movie 8. To show the resulting steady state, we measure the value of $\gamma$ where each transition occurs, and we plot the sequence of states as a function of time in Fig. 5D. The sequence matches that of Fig. 5B, it takes two periods to traverse, and it repeats indefinitely, as desired  [^5] There is one subtle difference in the pathways: The experiment navigates the downward avalanche via the unstable intermediate state +-+++-, whereas the design in Fig. 5A passes through ++-++- instead. We found that we could match the experiment by changing the bare threshold $\gamma_{1,0}^{-}$ for the top hysteron in Fig. 5A from 0 to $0.85$..
@@
Discussion. Recent work has demonstrated other mechanical designs that can count modulo two ten Wolde and Farhadi (2024), count driving cycles up to $n$ Kwakernaak and van Hecke (2023), or encode an analog mechanical input into a binary string Hyatt and Harne (2023). Although those systems use bistable mechanical elements, each one uses a different strategy to navigate the pathways between its states. In contrast, we have presented a general platform that can be reconfigured to perform all of the above functions, plus a latching behavior that was recently identified in hysteron simulations Lindeman et al. (2025). This platform could find use in the design of mechanisms that navigate through precise sequences of states, e.g., in robotics Kamp et al. (2024). More broadly, our work identifies a vast potential for designed materials that inherently engage with memory Keim et al. (2019), as their current state is a by-product of their driving history along with their overall design, which our platform now brings under the experimenter’s control.
@@
One principle of our approach was to use nonlinearities sparingly, so that the behaviors of the system could be rationalized as simply as possible. Nevertheless some important questions remain: Schrecengost Schrecengost (2025) showed that mixed interactions ($J_{ij}J_{ji}<0$) can be ruled out in a certain region of parameter space, but a full treatment of this question requires further work. Other designs for interacting mechanical hysterons are also being proposed; Shohat and van Hecke Shohat and van Hecke (2025) use theory and simulations to explore two-dimensional networks of bilinear hysterons as a route to exotic pathways, including mixed interactions and multiperiodic cycles. Understanding what are the general constraints on $\gamma_{i,0}^{\pm}$ and $J_{ij}$ for any realizable structure or metamaterial is an important open question Baconnier et al. (2025).
Designing the response of larger collections of mechanical hysterons presents another challenge Teunisse and van Hecke (2025); Muhaxheri et al. (2025). In Fig. 5 we obtained a period-2 response by aiming for a set of target hysteron parameters that were originally found via a random search Keim and Paulsen (2021). Our approach for dialing in that precise set of couplings was to iteratively adjust the mounting positions of springs and the post angles, to pick up more and more of the sought after multiperiodic cycle. This process is reminiscent of supervised learning Stern and Murugan (2023); Altman et al. (2024), suggesting that learning might be a viable approach for targeting transition graphs at large. Our system offers a wealth of possible learning degrees of freedom, as there are numerous geometric parameters that can be tuned continuously to affect the switching thresholds, including $x_{i}$, $\theta_{i}^{\pm}$, $\ell_{ij}$, and $L_{i}$. Further work is needed to determine whether a suitable learning algorithm could be used to navigate between different transition graphs, or classes of transition graphs with a desired functionality.
@@
Methods. Rotor hysteron construction. We construct each rotor hysterons from two 1/21/2”-diameter optical posts attached in a T configuration, which we mount in a radial steel ball bearing affixed to an optical table. We remove the bearing grease to further lower their friction to rotation. We use steel extension springs to couple the rotors to one another and to a 1/2”-diameter steel shaft that is mounted in two linear bearings. This 72”-long rod translates freely along the horizontal direction and provides the global mechanical drive, $\gamma$. We define $\gamma$ as the horizontal displacement of this rod in centimeters, which we measure with a precision of $0.1$ cm via a ruler affixed under the rod. We define $x_{i}$, the mounting position of driving spring $i$ on the driving rod, such that $x_{i}=0$ corresponds to no shift when the driving rod is at $\gamma=0$.